OpenStudy (mathlegend):

Write the equation of the line in slope-intercept form. perpendicular to y= 2x+6 and passing through (-4,2)

OpenStudy (mathlegend):

do I still use the point slope form.. formula?

OpenStudy (mathlegend):

y-y1=m(x-x1)

OpenStudy (anonymous):

yes that is correct and then you substitute the given values.

OpenStudy (mathlegend):

Do I keep the slope as 2?

OpenStudy (anonymous):

no, slope perpendicular to 2 is -1/2

OpenStudy (anonymous):

y = -1/2x

OpenStudy (mathlegend):

okay the reciprocal

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

perpendicular means, the slope for the other line is negative reciprocal

OpenStudy (mathlegend):

I got a weird answer... y = \[\frac{ -1 }{ 2}x\] + 0

OpenStudy (mathlegend):



OpenStudy (mathlegend):

I checked the answer key and it says it is wrong....

OpenStudy (anonymous):

plug the values in correctly into the point-slope form

OpenStudy (anonymous):

yup it must be y=-1/2x+4

OpenStudy (mathlegend):

Yeah that is the answer

OpenStudy (mathlegend):

What did you do that I did not do?

OpenStudy (mathlegend):

y-2=-1/2(x+4)

OpenStudy (anonymous):

\[y-2=m[x-(-4)]\] the perpendicular slope of 2 is -1/2

OpenStudy (mathlegend):

yeah I got that...

OpenStudy (anonymous):

i think that the answer in your notes is wrong and it should be y=-1/2x as y=-1/2(x+4)+2 simplifies to y=-1/2x-2+2 thus y=-1/2x

OpenStudy (anonymous):

\[y-2=-\frac{ 1 }{ 2 }[x-(-4)]\]

OpenStudy (mathlegend):

Oh so I was right... y=-1/2x + 0

OpenStudy (anonymous):

when you add 2 on both sides, what happens to the right side?

OpenStudy (anonymous):

it looks like it. The notes went wrong on the substitution. to get the answer that you have in the notes you need to have y=-1/2(x-4)+2

OpenStudy (anonymous):

you are correct: \[y-2+2=-\frac{ 1 }{ 2 }x-2+2\]